ane is an engineer with Southwest Airlines. She purchased Emmar stock for $6.90 per share and sold it exactly 1 year later for $13.14 per share. She was very pleased with her investment earnings. Help Jane understand exactly what she earned in terms of:

Effective annual rate=........%
Answer #1

The effective interest rate she earned = % change in the price of the stock = (13.14 - 6.90)/6.90 * 100 = 6.24/6.90 * 100 = 624/6.90 = 90.43%

Answer: 90.43%